OpenStudy (anonymous):

Monomials - Alg 2 (-4a^3b^2)^2*(3a^2b)

OpenStudy (jdoe0001):

\(\bf (-4a^3b^2)^2\times (3a^2b)\impliedby \textit{distribute the exponent} \\ \quad \\ (-4)^2(a^3)^2(b^2)^2\times 3a^2b^1\impliedby \textit{expand and combine same-bases} \\ \quad \\ (-4\cdot -4)(a^{3\cdot 2})(b^{2\cdot 2})\times 3a^2b^1 \\ \quad \\ 16a^6a^2b^4b^1\cdot 3\implies ?\)

OpenStudy (jdoe0001):

recall that \(\large a^n\cdot a^m = a^{n+m}\)

OpenStudy (anonymous):

I got 48a^8b^5 for my answer but it's probably wrong. I honestly don't understand monomials.

OpenStudy (jdoe0001):

\(\large \bf 16a^6a^2b^4b^1\cdot 3\implies 3\cdot 16a^{6+2}b^{4+1}\implies 48a^8b^5 \huge \checkmark\)

OpenStudy (anonymous):

Sweet. Thank you!

OpenStudy (anonymous):

Can you check this one?

OpenStudy (anonymous):

(-8x^4y^3) * (2x^5y^2) + 7x^9y^5

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\bf (-8x^4y^3) \cdot (2x^5y^2) + 7x^9y^5 \\ \quad \\ -8x^4y^32x^5y^2 + 7x^9y^5\implies -8x^4x^5y^32y^2 + 7x^9y^5 \\ \quad \\ -8x^{4+5}y^{3+2}+7x^9y^5\implies -8x^9y^5+7x^9y^5 \\ \quad \\ -8{\color{brown}{ x^9y^5 }} +7{\color{brown}{ x^9y^5 }} \implies -1{\color{brown}{ x^9y^5 }} \iff -x^9y^5 \)

OpenStudy (jdoe0001):

notice, the variables in the end, make up the "same group" or entity thus, you can add/subtrac them